OpenStudy (clara1223):

find the limit as x approaches 0 of (2x^2)/(tan^2)(9x)

OpenStudy (anonymous):

let me guess, no l'hopital?

OpenStudy (clara1223):

not yet

OpenStudy (thomas5267):

\[ \lim_{x\to0}\frac{2x^2}{\tan^2(9x)}? \]

OpenStudy (clara1223):

yes

OpenStudy (anonymous):

does it help to know that \[\lim_{x\to 0}\frac{x}{\tan(x)}=1\]?

OpenStudy (anonymous):

by which i really mean, can you use that?

OpenStudy (clara1223):

Yes we can use that, I'm just not sure how to get it to look like that

OpenStudy (anonymous):

algea bra pull out the 2, make a product out of it

OpenStudy (clara1223):

could I also pull out the 9?

OpenStudy (clara1223):

to make a product of 2/9?

OpenStudy (anonymous):

no the 9 is inside the tangent, but you can multiply top and bottom by 9

OpenStudy (anonymous):

or maybe you need 81 since there are two of them

OpenStudy (anonymous):

\[\frac{2}{81}\frac{9x}{\tan(9x)}\times \frac{9x}{\tan(9x)}\]

OpenStudy (clara1223):

so the answer is 2/81 x 1 x 1?

OpenStudy (anonymous):

yup

OpenStudy (thomas5267):

\[ \begin{align*} &\phantom{{}={}}\lim_{x\to0}\frac{2x^2}{\tan^2(9x)}\\ &=2\left(\lim_{x\to0}\frac{x}{\tan(9x)}\right)^2\\ &=2\left(\lim_{x\to0}\frac{x\cos(9x)}{\sin(9x)}\right)^2\\ &=2\left(\lim_{x\to0}\frac{x}{\sin(9x)}\right)^2\left(\lim_{x\to0}\cos(9x)\right)^2\\ &=2\left(\lim_{x\to0}\frac{1}{\frac{\sin(9x)}{x}}\right)^2\\ &=2\left(\frac{1}{\lim_{x\to0}\frac{\sin(9x)}{x}}\right)^2\\ &=2\left(\lim_{x\to0}\frac{\sin(9x)}{x}\right)^{-2}\\ &=2\left(\lim_{x\to0}\frac{9\sin(9x)}{9x}\right)^{-2}\\ &=2\left(9\lim_{x\to0}\frac{\sin(9x)}{9x}\right)^{-2}\\ &=2\left(9\right)^{-2}\\ &=\frac{2}{81} \end{align*} \]